# Simulations of Global Solar Convection with a Fully Compressible CHORUS++ Code

### v2

@@ 3.5 Spatial scale in the latitudinal and longitudinal directions
In panels (e) and (f), the ratio of the powers of $V_{\theta}$ to $V_{r}$ and $V_{\theta}$ to $V_{\phi}$ are shown. If the three components are evenly driven, then the ratio would be close to 1. As seen in these two panels, the power ratios are seldom equal or close to 1 ($10^{0}$); hence, the anisotropy in the convection zone (e.g., Guerrero et al., 2022) is significant. In the presented simulation, the radial component of plasma flow is set to zero ($V_{r}=0$) on the bottom and top boundary surfaces; hence, the ratios for depths close to the boundaries tend to be larger than the others.
@@
At the depths away from the boundary surfaces, overall, the ratio is greater than 1 for $l<10\sim 20$, implying that the horizontal flows dominate over the radial flows in the large scale over the wide range of depths. For larger harmonic degrees (i.e., $l>20$), the power ratios are rather constant at $10^{-0.3}\sim 10^{-0.4}$ or 0.4 $\sim$ 0.5 over the different depths; hence the square of the radial component $V_{r}^{2}$ and that of horizontal components $V^{2}_{h}=V_{\theta}^{2}+V_{\phi}^{2}$ are nearly balanced, indicating that the anisotropy of the plasma flows are small in the small spatial scale over a wide range of radius.
3.6 Long-term variations
Figure 10: Butterfly (latitude-time) diagram of the plasma quantities. The colors were truncated at the 95 percentile of the absolute values of each variable.
@@ 3.6 Long-term variations
Figure 11: Stack plot format of the plasma flow components. The colors were truncated at the 95 percentile of the absolute values of each variable.
@@
In Figure 11, the three components ($V_{r}$, $V_{\theta}$, and $V_{\phi}$) are plotted on the longitude-time diagram. The time here runs from top to bottom, following the convention of the stack plot format. In the top row, the simulated plasma flows at the equator (noted as ”lat=0” in the plot title). The prograde motions of the patterns can be clearly seen as the streaks running from top left to bottom right. This is straightforwardly related to the solar-like differential rotation shown in the meridional section view in Figure 6(f). In the middle row, the same properties but sampled at 45 degrees South are shown. The prograde motions are still noticeable. In the bottom row, the same quantities near the bottom boundary surface ($r=0.709\,R_{\sun}$) are shown. At this depth, no distinct tendency of the solar rotation can be found in part because of the boundary condition for the bottom boundary sphere.
4 Discussion and Summary
We conducted a hydrodynamic simulation of the solar convection zone with the CHORUS++ code. We used solar luminosity and solar rotation rate, and the results are substantially different from the benchmark tests by (Chen et al., 2023). Therefore, simulations with actual solar parameters are critically important to tackle many unanswered questions about the solar interior, solar dynamo, and solar-cycle variations (Fan, 2021, and references therein).
@@ 4 Discussion and Summary
This article offers the model descriptions and preliminary analysis of the simulation results. One item worth repeating here is the evaluation of the appropriateness of the anelastic and/or constant-mass assumptions that have been widely used. As shown in Section 3.4, the estimation least favoring the anelastic model still supports the divergence-free condition of mass flux and the sound-speed reduction method; in fact, the horizontal (latitudinal and longitudinal) gradients of the simulated mass density and temperature with the compressible HD code are very small with a deviation ratio of about $10^{-5}-10^{-4}$ relative to the average except for the region close to the solar surface. Therefore, we find that the divergence-free constant-mass density assumption is appropriate for studying the dynamics of the deep solar convection zone. At the same time, small gradients of the plasma properties are the primary drivers of the plasma motions; hence, the small-amplitude fluctuation of the plasma properties can affect the long-term dynamics. An implication of this study is that the contribution of the near-surface density fluctuations must be considered when the model includes the shallow subsurface region, as the time scale there is estimated at an order of day. A critically important expansion of the convection model is to set the upper boundary surface as close to the solar surface ($r=1\,R_{\sun}$) with much higher resolution (smaller grid size) in the radial direction to capture the plasma gradients near the solar surface and properly simulate the most turbulent part of the Sun.
@@
With a set of parameters, the solar-like differential rotation was obtained, although the meridional circulation patterns are not consistent with observations. In the presented model setting, the magnetic field is not included and the near-surface heat-escaping mechanism is simply expressed with the radiation dissipation term that is a function of radius. The results will improve or be more realistic if we include the magnetic field and MHD interactions in the convection zone (e.g., Hotta et al., 2022). The uniform heat flux is assumed at the bottom of the convection zone in the presented simulations. This assumption can be modified to examine the influence of the horizontal (latitudinal and longitudinal) gradients of the plasma quantities on the long-term solar dynamo (e.g., Bekki et al., 2024). In the presented study, we set the radial component of the plasma flow to zero $V_{r}=0$ on the boundary spheres together with the stress-free conditions for the horizontal components. In the actual Sun, the penetration of convection flows into the radiation zone is possible and can have a large influence on the entire dynamics of the solar/stellar cycles (e.g., Korre & Featherstone, 2024). The interactions between the convective zone and the radiation zone across the tachocline can be assessed by setting the radius of the bottom boundary surface below 0.7 $R_{\sun}$ with the CHORUS++ model.
@@
Recently, a new MHD version of the CHORUS++ model has been intensely tested (Paoli & Liang, 2025). We plan on modifying and expanding the model capabilities to realistically model the dynamics of the solar interior. This new version will allow investigations of the convection dynamics that have not yet been fully understood.
@@
This work was supported by the NASA DRIVE Center COFFIES (Consequences Of Fields and Flows in the Interior and Exterior of the Sun) grant 80NSSC22M0162. Computational resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. C.L. thanks a National Science Foundation (NSF) award (No. 2310372).

### v3

@@ III.5 Spatial scale in the latitudinal and longitudinal directions
In panels (e) and (f), the ratios of the powers of $V_{\theta}$ to $V_{r}$ and $V_{\theta}$ to $V_{\phi}$ are shown. If the three components are evenly driven then the ratio would be close to 1. As seen in these two panels, the power ratios are seldom equal or close to 1 ($10^{0}$); hence, the anisotropy in the convection zone (e.g., Guerrero et al., 2022) is significant. In the presented simulation, the radial component of plasma flow is set to zero ($V_{r}=0$) at the bottom and top boundary surfaces; hence, these ratios for depths closer to the boundaries tend to be larger.
@@
At the depths away from the boundary surfaces, overall, the ratio is greater than 1 for $l<10\sim 20$, implying that the horizontal flows dominate over the radial flows in the large scale over the wide range of depths. For larger harmonic degrees (i.e., $l>20$), the power ratios are rather constant at $10^{-0.3}\sim 10^{-0.4}$ or 0.4 $\sim$ 0.5 over the different depths; hence, the square of the radial component $V_{r}^{2}$ and that of horizontal components $V^{2}_{h}=V_{\theta}^{2}+V_{\phi}^{2}$ are nearly balanced, indicating that the anisotropy of the plasma flows is small in the small spatial scale over a wide range of radius.
Figure 10: Latitude-time diagram of the plasma velocity at a selected longitude (180 degrees). In the top row (a) – (c), the three components of the plasma velocity ($V_{r}$, $V_{\theta}$, $V_{\phi}$) sampled at $r=0.92R_{\sun}$ (near the top boundary surface) are shown. In the middle (d) – (f) and bottom row (g) – (i), the same quantities but sampled at $r=0.846R_{\sun}$ (the middle depth of the simulated domain) and at $r=0.709R_{\sun}$ (near the bottom boundary surface) are shown, respectively. The time runs from left to right. In panels (a) and (c), poleward motions of the patterns are found clearly. The colors were truncated at the 95th percentile of the absolute value of each variable.
Figure 11: The plasma velocity in the time-longitude plane (stack plot format). In the top row (a) – (c), the three components of the plasma velocity ($V_{r}$, $V_{\theta}$, $V_{\phi}$) sampled at $r=0.92R_{\sun}$ (near the top boundary surface) and on the equator (at the zero latitude) are shown. In the middle (d) – (f), the same quantities but sampled at 45 degrees south are shown. In the bottom (g) – (i), the velocity components sampled at $r=0.709R_{\sun}$ (near the bottom boundary surface) and on the equator are shown. The time runs from top to bottom, and the longitude runs from left to right. The prograde motions of the patterns are seen in the top row, and the slightly weaker but noticeable prograde motions are seen in the middle row. The colors are truncated at the 95th percentile of the absolute values of each variable.
III.6 Long-term variations
@@
In Figure 11, the three components ($V_{r}$, $V_{\theta}$, and $V_{\phi}$) are plotted on the longitude-time diagram. The time here runs from top to bottom, following the convention of the stack plot format. In the top row, the simulated plasma flows at the equator (noted as ”lat=0” in the plot title). The prograde motions of the patterns are clearly visible as the streaks running from top left to bottom right. This is straightforwardly related to the solar-like differential rotation shown in the meridional section view in Figure 6(f). In the middle row, the same properties, but sampled at 45 degrees South, are shown. The prograde motions are still noticeable. In the bottom row, the same quantities near the bottom boundary surface ($r=0.709\,R_{\sun}$) are shown. At this depth, no distinct tendency of the solar rotation can be found in part because of the boundary condition for the bottom boundary sphere.
IV Discussion and Summary
We conducted a hydrodynamic simulation of the solar convection zone with the CHORUS++ code. We used the values of solar luminosity and solar rotation rate, and the results are substantially different from the previous study (Chen et al., 2023). One of the differences is the growth of convection, as indicated by the temporal profile of the kinetic energy shown in Figure 3. In the previous study (Chen et al., 2023), the kinetic energy grows exponentially during the earliest phase, reaches a saturation state (or overshooting), then settles into a statistically stable equilibrium state. In the presented model parameters, although we also observed exponential growth of the kinetic energy, the transition from the saturated state to the stable state was very gradual and slow. Hence, this article analyzes the states that are still in transition.
Many of the parameters used in this study (tabulated in Table 1) are not equal to those estimated for the Sun. In particular, the parameters related to the fluid viscosity and the thermal conduction are set at orders of magnitude greater than the actual Sun. We could use the kinematic viscosity 10 times smaller than in (Chen et al., 2023); however, the value in the solar convection zone is much smaller by several orders of magnitude. Nonetheless, we here emphasize that the CHORUS++ code allows us to robustly conduct the simulations with relatively coarse grids. The efficiency of the code will allow us to conduct the simulations with the parameters much closer to the actual numbers of the solar convection zone, and better assess the dynamics of the solar interior, solar dynamo, and solar-cycle variations (Fan, 2021, and references therein).
@@ IV Discussion and Summary
This article offers the model descriptions and preliminary analysis of the simulation results. One item worth repeating here is the evaluation of the appropriateness of the anelastic models. As shown in Section III.4, the estimation least favoring the anelastic model still supports the divergence-free condition of mass flux and the sound-speed reduction method; in fact, the horizontal (latitudinal and longitudinal) gradients of the simulated mass density and temperature with the compressible HD code are very small with a deviation ratio of about $10^{-5}-10^{-4}$ relative to the average except for the region close to the solar surface. This result overall is consistent with a study comparing the fully compressible simulations and the anelastic models (Verhoeven et al., 2015) that found that the differences are overall small except for the regions with steep temperature gradients, like those at the outermost part of the solar convection zone.
Therefore, we find that the divergence-free constant-mass density assumption is appropriate for studying the dynamics of the deep solar convection zone. At the same time, small gradients of the plasma properties are the primary drivers of the plasma motions; hence, the small-amplitude fluctuation of the plasma properties can affect the long-term dynamics. An implication of this study is that the contribution of the near-surface density fluctuations must be considered when the model includes the shallow subsurface region, as the time scale there is estimated at an order of a day. A critically important expansion of the convection model is to set the upper boundary surface as close to the solar surface ($r=1\,R_{\sun}$) with much higher resolution (smaller grid size) in the radial direction to capture the plasma gradients near the solar surface and properly simulate the most turbulent part of the Sun.
@@
In the course of the transition of the convection system toward the equilibrium state, the simulated state exhibits a solar-like differential rotation pattern. The extension of the simulation time is necessary to confirm that the simulated system will eventually reach solar or anti-solar rotation patterns (e.g., Gilman, 1977; Gastine et al., 2014) and/or assess how long the presented patterns will last. In the presented model setting, the magnetic field is not included, and the near-surface heat-escaping mechanism is modeled as a radiation dissipation term that depends on radius. The results are expected to be more realistic when we include the magnetic field and MHD interactions in the convection zone (e.g., Hotta et al., 2022). The uniform heat flux is assumed at the bottom of the convection zone in the presented simulations. This assumption can be modified to examine the influence of the horizontal (latitudinal and longitudinal) gradients in plasma properties on the long-term solar dynamo (e.g., Bekki et al., 2024). In the presented study, we set the radial component of the plasma flow to zero, $V_{r}=0$, on the boundary spheres together with the stress-free conditions for the horizontal components. In the actual Sun, the penetration of convection flows into the radiation zone is possible and can strongly influence the entire dynamics of solar/stellar cycles (e.g., Korre and Featherstone, 2024). Furthermore, the physical processes acting in the radiative zone can substantially affect the tachocline and, hence, the convection zone. For example, Matilsky et al. (2022) conducted simulation studies for the region from 0.49 to 0.95 $R_{\sun}$ to assess the influences of the radiation zone on the solar cycle, and Brun et al. (2022) surveyed the influences in the sun-like stars with various values of the rotation rates and masses. In the future, the interactions between the convective zone and the radiation zone across the tachocline will be assessed by setting the radius of the bottom boundary surface below 0.7 $R_{\sun}$ with the CHORUS++ model.
@@
Recently, a new MHD version of the CHORUS++ model has been intensely tested (Paoli and Liang, 2025). We plan to modify and expand the model’s capabilities to capture the dynamics of the solar interior realistically. This new version will enable investigations of convection dynamics that are not yet fully understood.
@@
The authors appreciate the anonymous referee for constructive comments and suggestions, which helped improve this article. This work was supported by the NASA DRIVE Center COFFIES (Consequences Of Fields and Flows in the Interior and Exterior of the Sun) grant 80NSSC22M0162. Computational resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center. C.L. thanks a National Science Foundation (NSF) award (No. 2310372).